Is there a formula that could be used to explain how you add consecutive numbers to make multiples of 100.eg.18+19=19=20=21=22=100 38+39+40+41+42=200 that would be suitable for a twelve year old to be able to solve such problems?
Answer by checkley71(8403)

IF YOU DIVIDE 100 BY ANY ODD NUMBER (3,5,7,9,11,13,15,ETC) THAT GIVES YOU A WHOLE NUMBER (IE 100/5=20) THIS TELLS YOU THAT 5 SEQUENTIAL NUMBERS WITH 20 BEING THE MIDDLE NUMBER IS A SOLUTION: THUS 18+19+20+21+22=100
OR IF YOU DIVIDE 2600 BY 900/15=60 THUS 15 CONSECUTIVE NUMBERS WITH 60 AS THE
MIDDLE NUMBER WILL ADD UP TO 900 THUS
53+54+55+56+57+58+59+60+61+62+63+64+65+66+67=900
HOPE THIS HELPS. IT'S NOT VERY SIMPLE BUT IT DOES WORK.
TRY IT WITH 400/25=16 OR 25 NUMBERS WITH 16 AS THE MIDDLE NUMBER
TRY IT WITH 2200/11=200 OR 11 NUMBERS WIT 200 AS THE MIDDLE NUMBER